Joe's present age is 2/3 that of Gail. In 8 years, Joe's age will be 4/5 that of Gail's. How old are they now?
Answer by stanbon(75887)

Joe's present age is 2/3 that of Gail. In 8 years, Joe's age will be 4/5 that of Gail's. How old are they now?
--------------
Equatiions:
j = (2/3)g
j+8 = (4/5)(g+8)
-----
Substitute for "j" and solve for "g"::
(2/3)g + 8 = (4/5)g + (4/5)8
-----------------------
Multiply thru by 15 to get:
10g + 120 = 12g + 96
2g = 24
gail's age = 12 yrs (Gail's age now)
-----
Solve for "j":
j = (2/3)g
j = 8 yrs (Joe's age now)
===========
Cheers,
Stan H.
-----------